OpenStudy (anonymous):

Will give metal to most helpful!! Solve. Y-1=4y-2/3 y=_______

OpenStudy (dan815):

\[ \\ y-1=4y-\frac {2}{3} \\ y-1=\frac {4y-2}{3}\] Just to confirm, which one of those is your equation? The way you wrote it says is the first one

OpenStudy (anonymous):

it wants to know what y =equals

OpenStudy (anonymous):

the equation is y-1=4y-2/3, sorry i didnt understand what you meant at first

OpenStudy (dan815):

im just confirming* the 3 is only dividing 2?

OpenStudy (anonymous):

Your doing it right dan

OpenStudy (anonymous):

y-1=4y-2 ------- 3

OpenStudy (dan815):

okay i see you need to make sure you put brackets in properly, what you wrote doesnt say that

OpenStudy (anonymous):

oh, im sorry

OpenStudy (dan815):

its okay :)

OpenStudy (dan815):

now start by multiplying both sides by 3, what happens?

OpenStudy (dan815):

\[3*(y-1)=3*(\frac {4y-2}{3})\\\]

OpenStudy (dan815):

3*(Y-1)=3*(4y-2)/3

OpenStudy (dan815):

here is another way to see it

OpenStudy (anonymous):

3y-1=12y-2 -------- 3

OpenStudy (anonymous):

12y-6

OpenStudy (dan815):

you see now (12y-6)/3 =12y/3 - 6/3= 4y-2

OpenStudy (dan815):

also dont forget its 3y-3 on the left side

OpenStudy (dan815):